OpenStudy (anonymous):

Chris has $21.50. He wants a book for $5.85 and a CD for 14.99. The sales tax is 6.25% Does he have enough money to buy the book and the CD?

jimthompson5910 (jim_thompson5910):

Total (without tax) = 5.85 + 14.99 = 20.84 Now add 6.25% of that total onto the total 20.84 + 0.0625*20.84 = 22.1425 So the total with tax is $22.14 Since Chris only has $21.50, which is less than $22.14, this means he does NOT have enough money.

OpenStudy (gw2011):

No. $5.85+$14.99=$20.84 (20.84)(0.0625)=$1.30 tax $20.84+1.30=$22.14

OpenStudy (anonymous):

thanks both of you

OpenStudy (gw2011):

You're welcome